OpenStudy (anonymous):

The width of a square-based storage tank is 3m less than its height. The tank has a capacity of 20 m3. If the dimensions are integer values in metres, what are they?

OpenStudy (anonymous):

Let h denote the height, so the width is h-3. Volume = h(h-3)² = 20 h(h²-6h+9) = 20 h³-6h²+9h = 20 h³-6h²+9h-20 = 0 (h-5)(h²-h+4) = 0 h-5 = 0 h = 5 The dimensions are 5m by 2m by 2m. I hope this helps.

OpenStudy (anonymous):

how did you get from \[(h-5)(h ^{2}-h+4)=0\] to (h-5)=0 ?

OpenStudy (anonymous):

--->(h-5)<---(h²-h+4) =--->0<---=h-5 = 0

OpenStudy (anonymous):

i still don't get it...

OpenStudy (anonymous):

because (H-5)(h2-h+4)=0 so we put it together=H-5 so H-5=0 and that's the answer

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

I still domt get why h-5. Equals to zero . .